OpenStudy (cutestgirlaround):

please help!!! 3+17p=−32−11p slove for p

OpenStudy (qwertty123):

Hey, Let me see here...

OpenStudy (jiteshmeghwal9):

Collect all the like terms on one side

OpenStudy (qwertty123):

First you would... Simplify both sides of the equation. 3+17p=−32−11p

OpenStudy (qwertty123):

Then you would add 11p to both sides. 17p+3+11p=−11p−32+11p 28p+3= ?

OpenStudy (cutestgirlaround):

25p?

OpenStudy (qwertty123):

No...

OpenStudy (cutestgirlaround):

Oh :(

OpenStudy (qwertty123):

What is 28+3?

OpenStudy (cutestgirlaround):

oh 31 sorry I did 28 -3 lol

OpenStudy (qwertty123):

lol fine... but remember there is a negative... when you set it up... so it would be -32...

OpenStudy (qwertty123):

Then..Subtract 3 from both sides. 28p+3−3=−32−3 28p=−35

OpenStudy (cutestgirlaround):

then what's next

OpenStudy (qwertty123):

Then... Divide both sides by 28. To get your answer 28p over 28 = −35 over 28

OpenStudy (cutestgirlaround):

so its 28 over 28 = 1 the -35 = -1.25

OpenStudy (cutestgirlaround):

right?

OpenStudy (qwertty123):

no....

OpenStudy (cutestgirlaround):

I did -35 / 28 in my cal and its -1.25

OpenStudy (qwertty123):

No.. This you can't do on calculators...

OpenStudy (cutestgirlaround):

oh... then what do I do?

OpenStudy (qwertty123):

Um... hard to explain...@JFraser

OpenStudy (cutestgirlaround):

@Loli @andrejrusa12

OpenStudy (cutestgirlaround):

@Lolipop_Drum312

OpenStudy (cutestgirlaround):

I really need help

OpenStudy (lolipop_drum312):

i'm not really good with math but let me see it.. and i'll see what ic an do

OpenStudy (cutestgirlaround):

ok

OpenStudy (cutestgirlaround):

@T-Dawg02 you good at math?

OpenStudy (cutestgirlaround):

@AloneS

OpenStudy (cutestgirlaround):

@Unavailabilityy

alones (alones):

Alright were did you stop?

OpenStudy (lolipop_drum312):

hmm.... i might have worked it out 3+17p=-32-11p +11p +11p 3+28p =-32 -3 -3 28p = -32 you might wanna get comeone to check my answer though

OpenStudy (cutestgirlaround):

28 over 28 then -35 over 28 I think

alones (alones):

Hrmm i got something different /:

OpenStudy (lolipop_drum312):

ok spill it ^

OpenStudy (cutestgirlaround):

I know right I really would like starting over because I'm so confused

OpenStudy (lolipop_drum312):

i was just thinking the result could be -35/6

OpenStudy (t-dawg02):

p= -5/4

alones (alones):

^That's what I got..but I'm not sure Lol

alones (alones):

Maybe I'm wrong.

OpenStudy (lolipop_drum312):

it might be right

OpenStudy (lolipop_drum312):

hey cutes do you have an idea for this

OpenStudy (lolipop_drum312):

of what kinda answer you have to put in

OpenStudy (t-dawg02):

I'm 99.9999999% positive that it is p= -5/4

OpenStudy (cutestgirlaround):

all I know is I have to find out p what I normally would do is combine like terms in this situation so I would do 11p + 17p which = 28p then 3+ -32 which = 29 so I dont know

OpenStudy (cutestgirlaround):

I'm thinking @T-Dawg02 is right

OpenStudy (t-dawg02):

I am :D

OpenStudy (lolipop_drum312):

me too

OpenStudy (cutestgirlaround):

thanks everyone for your help i just dont know who to pick for the medal so can evryone just give each other a medal

OpenStudy (t-dawg02):

But i got it right